OpenStudy (anonymous):

Find the slope-intercept form of the given line that passes through the points (0, -3) and (4, 1). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (abmon98):

\[m=\Delta y/\Delta x=y _{2}-y _{1}/x _{2}-x _{1}\] \[(0,-3) (4,1) \]\[m=1-(-3)/4-0=1\] Slope-intercept form\[ y=mx+c \] or \[y-y1=m(x-x1) \]\[y-(-3)=1(x-0) \]\[y+3=x \]\[y=x-3 \]